OpenStudy (anonymous):

Determine whether the equation defines y as a linear function of x. If so, write it in the form y = mx + b. If not, enter NONE. 8x + 6y = 2 (a) Is this function linear?

OpenStudy (ghazi):

yes it is a linear function and it could be written as y=(2-8x)/6= (1-4x)/3

OpenStudy (ghazi):

for a function to be linear...highest order (or power or exponent) of term in the equation should be one

OpenStudy (anonymous):

typing it in atm

OpenStudy (ghazi):

y=(-4/3)x+(1/3)

OpenStudy (ghazi):

does it help you?

OpenStudy (anonymous):

May I define the term "linear function" in a slightly different statement:

OpenStudy (anonymous):

The dependence of Result-y upon the Variable-x is called linear when it is of the form \[y = m*x + b \] where m is A NUMBER , and called slope/coefficient and b is another number (sometimes called "intersection")

OpenStudy (anonymous):

So Z.W. - the answer of gh. is of the form that does indeed fit the definition of a linear func.

OpenStudy (anonymous):

Moreover -

OpenStudy (anonymous):

He also computed for u the actual numbers m and b

OpenStudy (anonymous):